Solve:y``-3y`+2y=e²t;y(0)=3;y-1(0)=5 by laplace transform


1
Expert's answer
2021-01-11T18:37:11-0500

Let us solve "y''-3y'+2y=e^{2t},\ y(0)=3,\ y'(0)=5" using laplace transform:


"y\to Y,\ \ y'\to pY-3,\ \ y''\to p^2Y-3p-5,\ \ e^{2t}\to \frac{1}{p-2}"


"p^2Y-3p-5-3pY+9+2Y=\frac{1}{p-2}"


"Y(p^2-3p+2)-3p+4=\frac{1}{p-2}"


"Y(p^2-3p+2)=\frac{1}{p-2}+3p-4=\frac{3p^2-10p+9}{p-2}"


"Y=\frac{3p^2-10p+9}{(p^2-3p+2)(p-2)}=\frac{3p^2-10p+9}{(p-1)(p-2)^2}"


"\frac{3p^2-10p+9}{(p-1)(p-2)^2}=\frac{A}{p-1}+\frac{B}{p-2}+\frac{C}{(p-2)^2}=\frac{A(p-2)^2+B(p-1)(p-2)+C(p-1)}{(p-1)(p-2)^2}"


"3p^2-10p+9=Ap^2-4Ap+4A+Bp^2-3Bp+2B+Cp-C"


"3p^2-10p+9=(A+B)p^2+(-4A-3B+C)p+4A+2B-C"


It follows that we have the following system:


"\begin{cases}
A+B=3\\
-4A-3B+C=-10\\
4A+2B-C=9
\end{cases}"


"\begin{cases}
A=2\\
B=1\\
C=1
\end{cases}"


Therefore, "Y=\frac{2}{p-1}+\frac{1}{p-2}+\frac{1}{(p-2)^2}"


Since "\frac{1}{p-1}\to e^t,\ \ \frac{1}{p-2}\to e^{2t},\ \ \frac{1}{(p-2)^2}\to te^{2t}," we conclude that the solution of the differential equation is the following:


"y(t)=2e^t+e^{2t}+te^{2t}."